OpenStudy (melissa_something):

Hello! :) Find the average rate of change in the function: f(x)=−3x2−x from x1=5 to x2=6

OpenStudy (mathmath333):

First find \(f(x_1)=f(5)\) and \(f(x_2)=f(6)\).

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} f(x)=−3x^2−x \hspace{.33em}\\~\\ f(5)=−3(5)^2−5=-75-5=-80 \hspace{.33em}\\~\\ \end{align}}\) similarly find \(f(6)\)

OpenStudy (melissa_something):

Oh thanks for correcting my math mistake. Oh okay and after that just subtract? @mathmath333

OpenStudy (mathmath333):

the formula would be \(\text{The average rate of change }=\dfrac{f(6)-f(5)}{6-5}\) so just find the value of \(f(5)\) and put it's value in the formula and get the answer.